OpenStudy (anonymous):

mike is making meatloaf he needs 7 pounds of meat. if he uses a 3:1ratio of beef to pork how much pork will he use?

OpenStudy (shane_b):

It's a ratio so: \[\frac{3}{1}=\frac{7}{x}\]Just solve for x.

OpenStudy (anonymous):

2.33?

OpenStudy (shane_b):

Yes!

OpenStudy (anonymous):

that was incorect..

OpenStudy (shane_b):

No it's not :) Maybe the system didn't like how you input the answer. Maybe try 2 1/3

OpenStudy (anonymous):

its 2 late i got it wrong lol

OpenStudy (shane_b):

:/

OpenStudy (shane_b):

Wait...I think I interpreted the question incorrectly. If he need 7 lbs of meat TOTAL, it would be:\[\frac{7}{4}=1.75\]

OpenStudy (shane_b):

I thought he was using 7 lbs of beef....my silly mistake.

OpenStudy (shane_b):

So his total recipe would be 5.25 lbs of beef and 1.75 lbs of pork.

OpenStudy (anonymous):

a lady went shopping and spent $54, there was a sale and every thing was 1/4 off what would she have payed without the sale?

OpenStudy (anonymous):

dammm i got 1.75!!!!!!

OpenStudy (shane_b):

Now I feel extra-bad :P

OpenStudy (anonymous):

dont work about it just help me with this new 1

OpenStudy (anonymous):

worry*

OpenStudy (shane_b):

0.75x=54

OpenStudy (shane_b):

Solving for x you'd get $72 as the non-sale price.

OpenStudy (anonymous):

ok that 1 was correct!

OpenStudy (shane_b):

Well, now I'm batting 50% :)

OpenStudy (anonymous):

brad buys 45$ of flowers daiseys are $1rosses are 2 he bought 3 more daiseys than roses how many roses did he buy?

OpenStudy (shane_b):

This is just a system of equations...which would look like this: x=roses, y= daisies\[1x+2y=45\]\[y=x+3\]Can you solve it from here?

OpenStudy (anonymous):

yes